find the length and width of a rectangle whose perimeter is 34 feet and whose area is 60 square feet.
Answer by ReadingBoosters(3246)

p = 34 = 2w+2l
a = 60 = wl; l = 60/w
...
2w + 2(60/w) = 34
2w + 120/w = 34
2w^2 + 120 = 34w
2w^2 - 34w + 120 = 0
(2w-10)(w-12)
w = 10/2 = 5
w = 12
...
lenth and width are 5ft and 12ft
...
Check
5*12 = 60
2(5) + 2(12) = 10 + 24 = 34

.....................
Delighted to help; feel free to email if you have further questions.
HomeworkHelpers@readingboosters.com
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.